OpenStudy (anonymous):

Please check my work for this one problem! X=0.2 & Y=3 when Z=2, & Z varies jointly with X & Y. z=kxy 2=k(0.2)(3) 2/0.6=k(0.6)/0.6 Z=3.3xy

OpenStudy (anonymous):

k=1/0.3,not 3.3.so z=xy/0.3

OpenStudy (anonymous):

Can you explain how you got that a little please. :b

OpenStudy (anonymous):

 alright?